# Focus on Your Question! Interpreting and Mitigating Toxic CoT Problems in Commonsense Reasoning

### v1

@@ Abstract
Focus on Your Question! Interpreting and Mitigating Toxic CoT Problems in Commonsense Reasoning
Jiachun Li 1,2, Pengfei Cao 1,2, Chenhao Wang 1,2, Zhuoran Jin 1,2, Yubo Chen 1,2 Daojian Zeng 3, Kang Liu 1,2, Jun Zhao 1,2 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Hunan Normal University {jiachun.li, pengfei.cao, chenhao.wang, zhuoran.jin, yubo.chen, kliu, jzhao} @nlpr.ia.ac.cn zengdj916@163.com
@@ 1 Introduction
With the increase in scale, large language models (LLMs) have demonstrated outstanding performance in different reasoning tasks, among them, commonsense reasoning has received significant attention due to its importance for general intelligence. In this task, researchers have proposed a series of chain-of-thought (CoT) like techniques to elicit models’ potential abilities (e.g. Self-Consistency (Wang et al., 2023c), Least-to-Most (Zhou et al., 2023), Reflexion (Shinn et al., 2023)). Through them, LLMs can generate reasonable rationales and improve their reasoning performance.
@@
While these works have made great progress, we notice an overlooked problem in them, which we define as Toxic CoT — Sometimes LLMs can directly provide correct answers to questions, but after applying CoT-like methods, it brings extra reasoning paths to models, causing their answers to be wrong. Figure 1 illustrates two main error types of this problem — Rationale Drift and Answer Drift. Specifically, for the Rationale Drift case, given the question “What kind of status is the bald eagle given?”, the model can directly give the correct answer “protection”. However, in the rationale, the model explains “what is the bald eagle” as “a symbol of America”, which has a semantic drift from the question. Thus, the model chooses the wrong option “america” based on the drifting rationale. For the Answer Drift case, given the question “Metal is used to make what?”, the model can directly answer “instruments”. It can also generate a correct rationale “metal is to make tools and machines”, but when answering based on the rationale, the model drifts from it and selects the incorrect option “(4)”. We further conduct a statistical analysis over extensive commonsense reasoning datasets and find that, among all CoT errors, this problem accounts for 37% for the white-box model and 33% for the black-box model on average, indicating this problem has become a crucial bottleneck in CoT reasoning. [^1] Appendix A presents the details of settings and results in this statistical experiment.
@@
(3) We conduct extensive experiments on various benchmarks. The results not only verify the rationality of our interpretation, but also demonstrate the effectiveness of our method in addressing the Toxic CoT problem (the proportion of the problem decreased by 23.6%) and enhancing commonsense reasoning performance (overall accuracy increased by 5.5%). We will release the source code for this work if it is accepted.
@@ 3 Tracing Information Flow in Rationale
[…]
@@ 4 Tracing Information Flow in Answer
[…]
@@ 4.2 Intervention Tracing Experiment
We report the result of Llama2-13B on Winogrande in Figure 5 and others in Appendix D, from which we can get two conclusions: (Claim 3) For attention modules, drifting cases loss information from the question. When the answer is correct, we can observe a high effect on the context and option in the first layer (see Figure 4(a)). But for the drifting case, the LLM extracts limited information at these positions (see Figure 4(c)). This aligns with Hypothesis 2. (Claim 4) For MLP modules, the information is not lost. We observe the same high-effect sites in the last layer and shallow layers of the last token, they do not show regular differences (see Figure 4(b) and 4(d)).
@@ 4.3 Attribution Tracing Experiment
[…]
@@ 5 Mitigating Toxic CoT Problem
[…]
@@ 5.3 Mitigation Results
The main result of our experiments on Llama2-13B is shown in Table 1 (see Table 8 for the result on Baichuan2-13B). We can get the following conclusions: (1) Our method effectively mitigates Toxic CoT problems. Compared to CoT prompting, our method reduces the Toxic Rate by an average of 23.6% across five datasets. Besides, compared to other advanced CoT-like methods, our method causes the fewest Toxic CoT problems (decreased by an average of 22.0% over SOTA methods). (2) Our method can also improve the model’s overall performance on commonsense reasoning. Our work improves the accuracy on all benchmarks (improved by 5.5% compared to CoT and 3.0% compared to SOTA methods on average). This proves that the Toxic CoT problem poses a bottleneck in LLM’s commonsense reasoning, highlighting the value of our work.
@@ 6.2 Mechanistic Interpretability
The work on mechanistic interpretability aims to understand the internal mechanisms of models when performing various tasks. Early work focused on how the model stores factual knowledge internally (Meng et al., 2022; Dai et al., 2022). In recent times, some research efforts have shifted towards examining how models retrieve and utilize knowledge. This includes internal knowledge retrieval (Geva et al., 2023), knowledge retrieval from prompts (Wang et al., 2023b), and the utilization of knowledge for reasoning purposes, such as math reasoning (Stolfo et al., 2023) and multi-step reasoning (Hou et al., 2023). However, there is limited existing work that explains commonsense reasoning and CoT reasoning, which are significant contributions of this work.
@@ 7 Conclusion
Although our work conducts an in-depth interpretation and mitigation of the Toxic CoT problem, it has several limitations. Firstly, like former commonsense reasoning works (Liu et al., 2022, 2023a; Xie et al., 2023), our research focuses on the form of multi-choice questions. This stems from the absence of effective evaluation methods for open-ended commonsense reasoning, leading to the predominance of benchmarks in this format. This calls for advancements in benchmark-related research. Secondly, we refrain from analyzing Toxic CoT problems in additional reasoning tasks such as math and logic, primarily due to the poor performance of current moderately-sized white-box models on these tasks. For instance, Llama2-13B achieves a mere 7.2% accuracy on GSM8K (Cobbe et al., 2021) without utilizing the CoT technique. This calls for developments in model-related research. We leave these limitations as our future work to explore.
@@ Appendix B Toxic Reason Statistical Experiments
[…]
@@ Appendix C More Details for Reasoning Tracing
[…]
@@
Strategy Wino CSQA Hella SIQA PIQA Avg
Greedy 5.2 2.8 4.9 3.6 5.8 4.5
Beam 8.1 4.9 10.1 5.8 9.4 7.7
Ours 13.4 6.6 16.2 7.4 15.9 11.9
Table 7: Decoding time per example comparison.
@@ Appendix F More Details for the Mitigation Experiment
For the Self-Consistency method, we sample 5 CoTs and use a majority voting method to select the final predicted answer. For the Self-Refine method, we first conduct one round of CoT reasoning and then follow it with one round of feedback to generate the final answer. For all the baselines, we will release the prompts in our source code if the work is accepted. We implement all of the methods on Llama2-13B-Chat-hf and Baichuan2-13B-Chat.
@@
Results
Table 8 reports more results in this experiment.
Appendix G More Details for Cost Analysis

### v2

@@ Abstract
Focus on Your Question! Interpreting and Mitigating Toxic CoT
Problems in Commonsense Reasoning
Jiachun Li1,2, Pengfei Cao1,2, Chenhao Wang1,2, Zhuoran Jin1,2, Yubo Chen1,2, [^1] Corresponding authors. Daojian Zeng3, Kang Liu1,2, Jun Zhao1,2, [^1] Corresponding authors. 1School of Artificial Intelligence, University of Chinese Academy of Sciences 2The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3Hunan Normal University {jiachun.li, pengfei.cao, chenhao.wang, zhuoran.jin, yubo.chen, kliu, jzhao} @nlpr.ia.ac.cn zengdj916@163.com
@@ 1 Introduction
With the increase in scale, large language models (LLMs) have demonstrated outstanding performance in different tasks (Li et al., 2023; Wen et al., 2023; Sun et al., 2024; Jin et al., 2024a), among them, commonsense reasoning has received significant attention due to its importance for general intelligence (Wang et al., 2022, 2024; Liu et al., 2024). In this task, researchers have proposed a series of chain-of-thought (CoT) like techniques to elicit models’ potential abilities (e.g. Self-Consistency (Wang et al., 2023c), Least-to-Most (Zhou et al., 2023), Reflexion (Shinn et al., 2023)). Through them, LLMs can generate reasonable rationales and improve their reasoning performance.
@@
While these works have made great progress, we notice an overlooked problem in them, which we define as Toxic CoT — Sometimes LLMs can directly provide correct answers to questions, but after applying CoT-like methods, it brings extra reasoning paths to models, causing their answers to be wrong. [^1] Notably, our definition here differs from the toxicity in generated content (Shaikh et al., 2023), emphasizing the harm that CoT can bring to the model. Figure 1 illustrates two main error types of this problem — Rationale Drift and Answer Drift. Specifically, for the Rationale Drift case, given the question “What kind of status is the bald eagle given?”, the model can directly give the correct answer “protection”. However, in the rationale, the model explains “what is the bald eagle” as “a symbol of America”, which has a semantic drift from the question. Thus, the model chooses the wrong option “america” based on the drifting rationale. For the Answer Drift case, given the question “Metal is used to make what?”, the model can directly answer “instruments”. It can also generate a correct rationale “metal is to make tools and machines”, but when answering based on the rationale, the model drifts from it and selects the incorrect option “(4)”. We further conduct a statistical analysis over extensive commonsense reasoning datasets and find that, among all CoT errors, this problem accounts for 37% for the white-box model and 33% for the black-box model on average, indicating this problem has become a crucial bottleneck in CoT reasoning. [^2] Appendix A presents the details of settings and results in this statistical experiment.
@@
(3) We conduct extensive experiments on various benchmarks. The results not only verify the rationality of our interpretation, but also demonstrate the effectiveness of our method in addressing the Toxic CoT problem (the proportion of the problem decreased by 23.6%) and enhancing commonsense reasoning performance (overall accuracy increased by 5.5%). Our code is available at: https://github.com/BugMakerzzz/toxic_cot.
@@ 3 Tracing Information Flow in Rationale
In this section, we aim to verify the Hypothesis 1 by tracing the information flow in the rationale generation stage. To this end, we start by describing our attribution tracing method (§$\S$3.1). Through this method, we conduct comparative experiments between the correct reasoning and the drifting one, figuring out the mechanism behind the issue (§$\S$3.2). At last, we use the attention score to validate our findings from another perspective (§$\S$3.3).
@@ 4 Tracing Information Flow in Answer
In this section, our goal is to verify the information loss based on Hypothesis 2. To achieve this, we first introduce the main tracing method in this section, which is called the causal tracing method (§$\S$4.1). Next, by employing it, we trace the information flow in the answer generation stage and identify the mechanism behind the Answer Drift issue through comparative experiments (§$\S$4.2). At last, we apply the attribution tracing method to verify our hypothesis from another perspective (§$\S$4.3).
@@ 4.2 Intervention Tracing Experiment
We report the result of Llama2-13B on Winogrande in Figure 5 and others in Appendix D, from which we can get two conclusions: (Claim 3) For attention modules, drifting cases loss information from the question. When the answer is correct, we can observe a high effect on the context and option in the first layer (see Figure 5(a)). But for the drifting case, the LLM extracts limited information at these positions (see Figure 5(c)). This aligns with Hypothesis 2. (Claim 4) For MLP modules, the information is not lost. We observe the same high-effect sites in the last layer and shallow layers of the last token, they do not show regular differences (see Figure 5(b) and 5(d)).
@@ 4.3 Attribution Tracing Experiment
For further validation of our hypothesis, we also use the attribution score in §$\S$3 to trace the information flow in this stage. Referring to Equation 2, we compute the score between the question context and the last token (since it’s used for generating the answer). We set the $F(\cdot)$ in Equation 1 as the loss for predicting the final answer, comparing the scores for correct and drifting cases after averaging across samples.
@@ 5 Mitigating Toxic CoT Problem
In this section, we propose a novel method called $\mathbb{RIDERS}$ (Residual decodIng and sERial-position Swap) to address the Toxic CoT problem. We first introduce the two components in it, which are designed based on Hypothesis 1 and 2, respectively (§$\S$5.1). Then, we conduct experiments on commonsense reasoning benchmarks, demonstrating the effectiveness of our method (§$\S$5.2 and §$\S$5.3). At last, we conduct extra experiments to further emphasize the contribution of our approach (§$\S$5.4).
@@ 5.3 Mitigation Results
The main result of our experiments on Llama2-13B is shown in Table 1. (The results on more models are presented in Appendix F.) We can get the following conclusions: (1) Our method effectively mitigates Toxic CoT problems. Compared to CoT prompting, our method reduces the Toxic Rate by an average of 23.6% across five datasets. Besides, compared to other advanced CoT-like methods, our method causes the fewest Toxic CoT problems (decreased by an average of 22.0% over SOTA methods). (2) Our method can also improve the model’s overall performance on commonsense reasoning. Our work improves the accuracy on all benchmarks (improved by 5.5% compared to CoT and 3.0% compared to SOTA methods on average). This proves that the Toxic CoT problem poses a bottleneck in LLM’s commonsense reasoning, highlighting the value of our work.
@@ 5.4 Discussion and Analysis
In §$\S$3 and §$\S$4, we probe information loss in two issues by tracing the information flow in models. Here, we repeat the attribution tracing experiments, comparing the differences before and after applying our method to further validate the effectiveness of our work. As we can see from Figure 8 and 16, our two methods (orange lines) increase the information flow from questions in two stages compared to CoT prompting (blue lines). This indicates our method indeed compensates for the information loss in the LLM.
Method ProofWriter FOLIO
@@
CoT 56.8 19.9 45.1 11.6
RIDERS 78.2 6.4 48.0 7.5
Table 3: Performance comparison across two logical reasoning tasks on Llama2-13B.
Performance on Other Tasks
Our work mainly focuses on commonsense reasoning tasks. To further evaluate the effectiveness of our approach on other types and forms of tasks, we conduct the main experiments on two logical reasoning tasks: ProofWriter (Tafjord et al., 2021) and FOLIO (Han et al., 2022)). As presented in Table 3, we can find that the Toxic CoT problem also exists in other tasks, and our method can still mitigate it.
@@ 6.2 Mechanistic Interpretability
The work on mechanistic interpretability aims to understand the internal mechanisms of models when performing various tasks. Early work focused on how the model stores factual knowledge internally (Meng et al., 2022; Dai et al., 2022). In recent times, some research efforts have shifted towards examining how models retrieve and utilize knowledge. This includes internal knowledge retrieval (Geva et al., 2023), knowledge retrieval from prompts (Wang et al., 2023b; Jin et al., 2024b), and the utilization of knowledge for reasoning purposes, such as math reasoning (Stolfo et al., 2023) and multi-step reasoning (Hou et al., 2023). However, there is limited existing work that explains commonsense reasoning and CoT reasoning, which are significant contributions of this work.
@@ 7 Conclusion
Although our work conducts an in-depth interpretation and mitigation of the Toxic CoT problem, it has several limitations. Firstly, like former commonsense reasoning works (Liu et al., 2022, 2023a; Xie et al., 2023), our research focuses on the form of multi-choice questions. This stems from the absence of effective evaluation methods for open-ended commonsense reasoning, leading to the predominance of benchmarks in this format. This calls for advancements in benchmark-related research. Secondly, we refrain from analyzing Toxic CoT problems in more reasoning tasks such as math, primarily due to the poor performance of current moderately-sized white-box models on these tasks. For instance, Llama2-13B achieves a mere 7.2% accuracy on GSM8K (Cobbe et al., 2021) without utilizing the CoT technique. This calls for developments in model-related research. We leave these limitations as our future work to explore.
Acknowledgement
This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDA27020203), the National Natural Science Foundation of China (No. 62176257,62276095).
@@ Appendix B Toxic Reason Statistical Experiments
As the inconsistency error constitutes the predominant portion of all reasons, our work focuses on addressing this issue. We further categorize this error into Rationale Drift and Answer Drift based on the error occurrence (see §$\S$2.2 for their definitions).
@@ Appendix C More Details for Reasoning Tracing
We introduce the attribution score method in §$\S$ 3.1. In Equation 1, we set $m$ = 20 following the previous works. For $F(\cdot)$, we set it as the language modeling loss (for next-token prediction) during the CoT generation. Here, we obtain this value directly from the output of the LlamaForCausalLM module using the Transformers library. In Equation 2, we partition the step numbers $|N|$ in CoT based on the occurrence of periods in the text. For models, we use Llama2-13B-Chat and Baichuan2-13B-Chat.
@@ Appendix F More Details for the Mitigation Experiment
For the Self-Consistency method, we sample 5 CoTs and use a majority voting method to select the final predicted answer. For the Self-Refine method, we first conduct one round of CoT reasoning and then follow it with one round of feedback to generate the final answer. For all the baselines, we release the prompts in our source code. We implement all of the methods on Llama2-13B-Chat-hf and Baichuan2-13B-Chat.
@@
Model Method Winogrande SIQA
[…]
Mistral-7B CoT 59.2 27.9 70.2 28.2
RIDERS 62.0 15.3 70.6 24.5
GPT-3.5 CoT 64.6 48.8 70.6 30.3
RIDERS 70.1 41.7 72.1 29.0
Table 9: Performance comparison on Mistral-7B and GPT-3.5.
Results
In our paper, we choose Llama2-13B for experiments because it is an open-source LLM of suitable size and has great influence. Here, we also repeat all experiments on Baichuan2-13B to verify the generality of our work (see Table 8). Additionally, in Table 9, we show our experimental results on two other representative models: Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-1106 (since it is a closed source model, here we only apply the SPS method).
Strategy Wino CSQA Hella SIQA PIQA Avg
Greedy 5.2 2.8 4.9 3.6 5.8 4.5
Beam 8.1 4.9 10.1 5.8 9.4 7.7
Ours 13.4 6.6 16.2 7.4 15.9 11.9
Table 10: Decoding time per example comparison.
Appendix G More Details for Cost Analysis